If {9}^{-1 + 2\,x} == \frac{{81}^{2+x}}{{3}^{x}} find x


1
Expert's answer
2021-11-08T20:27:42-0500

"{9}^{-1 + 2\,x} = \frac{{81}^{2+x}}{{3}^{x}}"

"(3^2)^{-1 + 2\,x} = (3^4)^{2+x}\cdot3^{-x}"

"3^{-2+4x}=3^{8+4x-x}"

"-2+4x=8+3x"

"x=10"